# Exploring the Frontier of Vision-Language Models: A Survey of Current Methodologies and Future Directions

### v1

@@ 2.2 Text generation with Multimodal Input
LLaVa-1.5 Liu et al. (2023a): It is a refined version of the LLaVA, focusing on visual instruction tuning to enhance multimodal models. The paper outlines modifications to LLaVA, such as using CLIP-ViT-L-336px with an MLP projection and incorporating academic-task-oriented Visual Question Answering (VQA) data. Despite its advancements, limitations are acknowledged, such as prolonged training iterations due to the use of full image patches and challenges in processing multiple images and certain domain-specific tasks.
@@
CogVLM Wang et al. (2023a): CogVLM is an open-source vision-language foundation model developed by researchers from Tsinghua University. Its architecture comprises a Vision Transformer (ViT) encoder (e.g., EVA2-CLIP-E) for image processing, with output mapped into the text feature space using an MLP adapter. The model includes a pre-trained GPT-style language model and a visual expert module added to each layer, consisting of a QKV matrix and an MLP. CogVLM adopts a deep fusion approach, integrating visual and language features at multiple layers through the visual expert module, surpassing traditional shallow alignment methods. Alignment techniques involve pretraining on a vast dataset of 1.5 billion image-text pairs, employing image captioning loss and Referring Expression Comprehension (REC). Fine-tuning on various tasks, with a focus on free-form instructions, leads to the creation of a variant known as CogVLM-Chat.
@@
FERRET You et al. (2023): FERRET is designed for spatial referring and grounding in images at different shapes and granularities. Ferret’s distinct features include a Hybrid Region Representation, blending discrete coordinates and continuous visual features for diverse region inputs. It uses a Spatial-Aware Visual Sampler to handle various region shapes effectively and is trained on the Ground-and-Refer Instruction-Tuning (GRIT) dataset, which includes hierarchical spatial knowledge and hard negative samples. The architecture includes an image encoder, a spatial-aware visual sampler, and a Language Model . Ferret utilizes a pre-trained visual encoder (CLIP-ViT-L/14) and a Language Model’s tokenizer for image and text embeddings. Training occurs on the GRIT dataset for three epochs, with the model randomly choosing between center points or bounding boxes to represent regions. In multimodal chatting tasks, Ferret significantly enhances performance by integrating refer-and-ground capabilities. Notably, Ferret mitigates the issue of object hallucination, a common challenge in multimodal models.
@@
BARDGoogle (2023): BARD from Google utilizes a reinforcement learning framework to automate machine learning model design, architecture search, and hyperparameter tuning, making it accessible to users without extensive AI expertise. The system is positioned as a standalone experiment, with a focus on productivity, creativity, and curiosity enhancement. Users engage BARD for tasks such as writing resumes, creating workout routines, and planning itineraries. The model is pre-trained on diverse data sources, and responses are generated by considering context, classified against safety parameters, and re-ranked based on quality. Human feedback and evaluation, including fine-tuning and reinforcement learning on human feedback, are used to improve BARD. Limitations include potential inaccuracies, biases, persona attribution, false positives/negatives, and vulnerability to adversarial prompting. Google is committed to addressing these limitations and improving BARD responsibly over time.
@@
LLaMA-VID Li et al. (2023c): LLaMA-VID introduces a novel dual-token strategy, incorporating context and content tokens, to efficiently encode each video frame. This approach enables the model to handle hour-long videos while mitigating computational complexity. LLaMA-VID employs a hybrid architecture, incorporating pre-trained models like Vicuna for text processing and a Vision Transformer for image embeddings in videos. The Q-Former introduces the context-attention token (Et) by computing attention between query-generated text embeddings (Q) and visual tokens (X). Et encapsulates relevant visual features. A content token (Ev) is obtained through mean pooling on visual tokens. Both tokens are integrated into the V decoder for generating text responses. LLaMA-VID’s dual-token generation strategy, comprising context and content tokens, ensures adaptability to various settings, optimizing efficiency for videos while preserving detail for single images.LLaMA-VID is a video and image understanding model designed for efficiency, completing training in two days on 8xA100 GPUs. It uses EVA-G for visual encoding and QFormer for text decoding. The training set includes image and video caption pairs, with evaluations on diverse benchmarks. LLaMA-VID excels in zero-shot video QA benchmarks, achieving high accuracy with only two tokens per frame.
@@
CoVLM Li et al. (2023b): CoVLM introduces a novel approach to enhance large language models’ compositional reasoning by integrating vision-language communicative decoding. Utilizing communication tokens, the model dynamically composes visual entities and relationships, improving language generation through iterative communication with vision encoders and detection networks . Trained on a large-scale dataset , CoVLM outperforms baseline vision-language models in compositional reasoning tasks, and exhibits competitive performance in tasks such as referring expression comprehension and visual question answering. The model represents a noteworthy advancement in integrating vision and language models, with acknowledgment of potential future improvements in compositionality.
@@
Emu2: Sun et al. (2023a): Emu2 is a 37 billion-parameter generative multimodal model, that achieves remarkable in-context learning on diverse multimodal sequences, setting new records in few-shot understanding tasks. Employing a unified autoregressive objective, Emu2 seamlessly integrates visual embeddings and textual tokens. Its architecture includes a visual encoder, multimodal modeling, and visual decoder, allowing coherent outputs across different modalities. Emu2 excels in vision-language tasks, instruction tuning, and controllable visual generation, showcasing state-of-the-art performance in image question answering, subject-driven generation, and zero-shot text-to-image generation. The paper acknowledges broader impact considerations and limitations, emphasizing responsible deployment in light of challenges such as hallucination, biases, and question-answering capabilities.
@@
Video-LLaMA Zhang et al. (2023c): Video-LLaMA is designed to understand both visual and auditory content in videos. The model integrates pre-trained visual and audio encoders with frozen LLMs, addressing challenges in capturing temporal changes in visual scenes and integrating audio-visual signals. Utilizing a Video Q-former for temporal information and an Audio Q-former for audio encoding, the framework aligns audio-visual data with textual information. Experimental results demonstrate Video-LaMA’s effectiveness in comprehending video content and generating meaningful responses in audio and video-grounded conversations. However, the paper acknowledges limitations such as restricted perception capacities and challenges with long videos. Despite these, Video-LLaMA represents a notable advancement in audio-visual AI assistants, with the authors providing open-sourced resources for further development.
@@
Video-ChatGPT Maaz et al. (2023): It is, a novel multimodal model enhancing video understanding by integrating a video-adapted visual encoder with a Large Language Model. The architecture leverages the CLIP ViT-L/14 visual encoder for spatiotemporal video representations and the V-v1.1 language model for comprehensive understanding. Notably, a dataset of 100,000 video-instruction pairs is created to fine-tune the model, focusing on temporal relationships and contextual understanding. The model exhibits competitive performance in correctness, detail orientation, contextual and temporal understanding, and consistency, surpassing contemporary models in zero-shot question-answering tasks. Qualitatively, Video-ChatGPT demonstrates proficiency in various video-based tasks but faces challenges in subtle temporal relationships and small visual details, indicating avenues for future improvement.
@@
LAVIN Luo et al. (2023): LAVIN is utilizing Mixture-of-Modality Adaptation (MMA) for cost-effective adaptation of LLMs to vision-language tasks. LaVIN, with lightweight adapters, achieves competitive performance and superior training efficiency in multimodal tasks like science question answering and dialogue. Remarkably, LaVIN requires only 1.4 training hours and 3.8M trainable parameters. Experimental results on the ScienceQA dataset demonstrate efficiency with comparable performance and reduced training time and storage costs. LAVIN represents a breakthrough in cost-effective adaptation but has limitations, including the potential for incorrect responses and challenges in identifying fine-grained details in images.
@@
BEiT-3 Wang et al. (2022): BEiT-3 is, a multimodal foundation model showcasing a significant convergence in language, vision, and multimodal pretraining. BEiT-3 excels in transfer performance for both vision and vision-language tasks, advancing convergence through innovations in backbone architecture, pretraining tasks, and scalable model design. Employing Multiway Transformers, the model features a modular architecture enabling deep fusion and modality-specific encoding. Utilizing a shared backbone, BEiT-3 conducts unified masked "language" modeling on images, texts (English), and image-text pairs ("parallel sentences"). Experimental results demonstrate BEIT-3 achieving state-of-the-art performance across various tasks, including object detection, semantic segmentation, image classification, visual reasoning, visual question answering, image captioning, and cross-modal retrieval.
@@
mPLUG-2 Xu et al. (2023a): Departing from prevalent approaches reliant on sequence-to-sequence generation or encoder-centric instance discrimination, mPLUG-2 introduces a multi-module composition network. This innovative design fosters modality collaboration while effectively addressing modality entanglement. The flexibility inherent in mPLUG-2 allows for the selective use of diverse modules across text, image, and video modalities for various understanding and generation tasks. Empirical evaluations demonstrate mPLUG-2’s prowess, achieving state-of-the-art or competitive results across an extensive range of over 30 downstream tasks. From challenging multi-modal endeavors like image-text and video-text understanding to uni-modal tasks spanning text-only, image-only, and video-only domains, mPLUG-2 exhibits its versatility. Particularly noteworthy is its achievement of new state-of-the-art performance with a 48.0 top-1 accuracy and 80.3 CIDEr on video QA and video caption tasks, all achieved with a notably smaller model size and data scale. Additionally, mPLUG-2 showcases robust zero-shot transferability in vision-language and video-language tasks, solidifying its position at the forefront of multimodal pretraining advancements.
@@
X2-VLM Zeng et al. (2023): X2-VLM is a versatile model with a flexible modular architecture that integrates image-text and video-text pre-training into a unified framework. It excels in image-text and video-text tasks across different scales, balancing performance and model scale. X2-VLM’s modular design enhances transferability, allowing seamless use in various languages or domains. By substituting the text encoder, it outperforms state-of-the-art multilingual multimodal pre-trained models, demonstrating superior performance without requiring specific multilingual pre-training. This adaptability positions X2-VLM as a promising model in the field of multimodal pre-training.
Lyrics Lu et al. (2023): Lyrics, an innovative paradigm in multi-modal pre-training and instruction fine-tuning that revolutionizes vision-language alignment through fine-grained cross-modal collaboration. Evolving from the robust foundation of BLIP-2, Lyrics incorporates local visual features extracted from a sophisticated visual refiner, encompassing image tagging, object detection, and semantic segmentation modules. In the Querying Transformer, these visual features seamlessly integrate with language inputs, enriched by boundary boxes and tags derived from the visual refiner. The two-stage training approach is a key highlight, addressing the modality gap by establishing explicit and comprehensive vision-language alignment targets during pre-training. Subsequently, in the instruction fine-tuning stage, we introduce semantic-aware visual feature extraction, a pivotal method empowering the model to glean informative features from tangible visual objects. The efficacy of the approach is underscored by its robust performance across 13 held-out datasets spanning diverse vision-language tasks.
@@
X-FM Zhang et al. (2023b): XFM, a novel general foundation model equipped with one language encoder, one vision encoder, and one fusion encoder, featuring a unique training method. The proposed method incorporates two innovative techniques: halting gradients from vision-language training during language-encoder learning and leveraging vision-language training to guide vision-encoder learning. Extensive experiments on benchmark datasets demonstrate that X-FM outperforms existing general foundation models and performs competitively with or surpasses models tailored specifically for language, vision, or vision-language understanding. The paper acknowledges limitations, including substantial computational requirements, and aims to explore techniques for efficiency improvement and reduced environmental impact. The authors highlight their commitment to addressing efficiency challenges and reducing the carbon footprint in line with "green" deep learning initiatives. However, due to computational constraints, the study did not explore super-large models or pre-train large-sized models on extensive datasets, emphasizing scalability as an essential consideration for foundation models.
@@
VALOR Chen et al. (2023a): VALOR is a unified vision-audio-language crossmodality pretraining model designed for trimodality understanding and generation. VALOR employs two pretraining tasks, Multimodal Grouping Alignment and Multimodal Grouping Captioning, showcasing good versatility and scalability. Two datasets, VALOR-1M and VALOR-32K, are proposed for tri-modality pretraining research and benchmarking audiovisual-language retrieval and captioning. Trained on VALOR-1M and other vision-language datasets, VALOR achieves new state-of-the-art performances in downstream tasks like vision/audio/audiovisual retrieval, captioning, and question answering. The paper outlines plans for future work, including scaling the VALOR-1M dataset through unsupervised methods and introducing vision and audio generation modeling into the VALOR framework.
@@
Prismer Liu et al. (2023b): Prismer is a data- and parameter-efficient vision-language model that utilizes a frozen ensemble of domain experts, minimizing the need for extensive training data. By inheriting weights from pre-trained domain experts across various domains and keeping them frozen during training, Prismer efficiently adapts to different vision-language reasoning tasks. Despite its small-scale language model foundation, Prismer demonstrates competitive fine-tuned and few-shot learning performance, requiring significantly less training data than current state-of-the-art models. However, it lacks the ability for zero-shot in-context generalization and shows limitations in adapting to new experts or partial expert ensembles during inference, leading to performance drops. The paper discusses these limitations, including the absence of few-shot in-context prompting, challenges in adapting to new experts, and potential improvements in representing expert knowledge for enhanced reasoning performance in future iterations.
@@
MMReact Yang et al. (2023b): MM-REACT introduces a novel textual prompt design enabling language models to process multimodal information, including text descriptions, spatial coordinates, and file names for dense visual signals. The approach demonstrates its effectiveness in zero-shot experiments, showcasing its potential for advanced visual understanding across various scenarios. However, the paper identifies limitations, such as challenges in systematically evaluating performance due to the absence of annotated benchmarks for recognition capability in the wild. The integrated vision experts may introduce errors, and the system’s success depends on the availability of necessary experts. Additionally, the number of experts is constrained by the context window of ChatGPT, and the conversion of visual signals to text words may not be optimal for certain tasks. Manual prompt engineering is required, and the authors suggest future research to automate this process for increased system development ease.
@@
PICa Jin et al. (2021): PICa is a method utilizing image captions to prompt GPT-3 for knowledge-based Visual Question Answering (VQA). Leveraging GPT-3’s knowledge retrieval and question-answering capabilities, the approach treats GPT-3 as an implicit and unstructured knowledge base, converting images into captions or tags for GPT-3 understanding. By adapting GPT-3 for VQA through a few-shot learning approach with in-context examples, PICa achieves notable performance, surpassing the supervised state of the art on the OK-VQA dataset with just 16 examples. The method is the first to use GPT-3 for multimodal tasks. However, a limitation is noted, as the image is abstracted as text, and captions may provide only a partial description, potentially missing crucial visual details necessary for detailed question answering, such as queries on specific visual attributes.
@@
PNP-VQA Tiong et al. (2022): Plug-and-Play VQA (PNP-VQA) is a modular framework designed for zero-shot Visual Question Answering (VQA). Unlike existing approaches that demand extensive adaptation of pre-trained language models (PLMs) for vision, PNP-VQA eliminates the need for additional training of PLMs. Instead, it employs natural language and network interpretation as an intermediate representation to connect pre-trained models. The framework generates question-guided informative image captions, utilizing them as context for PLMs during question answering. PNP-VQA surpasses end-to-end trained baselines and achieves state-of-the-art results on zero-shot VQAv2 and GQA datasets. With 11B parameters, it outperforms an 80B-parameter model on VQAv2 and exhibits a 9.1% improvement over a comparable model on GQA, showcasing its efficacy with different PLM parameter sizes.
@@
Img2LLM Guo et al. (2023): Img2LLM is designed for LLMs that facilitates zero-shot VQA without necessitating end-to-end training. The approach involves developing LLM-agnostic models that articulate image content through exemplar question-answer pairs, proving to be effective prompts for LLMs. Img2LLM boasts several advantages, achieving performance on par with or surpassing end-to-end trained methods, such as outperforming Flamingo by 5.6% on VQAv2 and exhibiting notable superiority on the challenging A-OKVQA dataset. Additionally, the flexibility of Img2LLM allows seamless integration with various LLMs for VQA tasks, eliminating the need for specialized, costly end-to-end fine-tuning. One caveat is the additional inference overhead incurred during image caption and question-answer pair generation, contributing to a 24.4% increase in computational time. However, this overhead can be mitigated by shortening prompts, trading a fraction of accuracy for speed, while Img2LLM avoids the resource-intensive end-to-end multimodal representation alignment seen in comparable models like Flamingo.
@@
SimVLM Wang et al. (2021): SimVLM is a streamlined pretraining framework that embraces a minimalist approach. Unlike previous methods, SimVLM simplifies training complexities by leveraging large-scale weak supervision, and undergoing end-to-end training with a singular prefix language modeling objective. Remarkably, without resorting to additional data or task-specific tailoring, the resultant model surpasses its predecessors like OSCAR, VILLA etc establishing new benchmarks in various vision-language tasks. Additionally, SimVLM demonstrates robust generalization and transfer capabilities, showcasing zero-shot behavior in tasks such as open-ended visual question answering and cross-modality transfer.

### v2

@@ 2.2 Text generation with Multimodal Input
LLaVa-1.5 [51]: It is a refined version of the LLaVA, focusing on visual instruction tuning to enhance multimodal models. The paper outlines modifications to LLaVA, such as using CLIP-ViT-L-336px with an MLP projection and incorporating academic-task-oriented Visual Question Answering (VQA) data. Despite its advancements, limitations are acknowledged, such as prolonged training iterations due to the use of full image patches and challenges in processing multiple images and certain domain-specific tasks.
@@
CogVLM [85]: CogVLM is an open-source vision-language foundation model developed by researchers from Tsinghua University. Its architecture comprises a Vision Transformer (ViT) encoder (e.g., EVA2-CLIP-E) for image processing, with output mapped into the text feature space using an MLP adapter. The model includes a pre-trained GPT-style language model and a visual expert module added to each layer, consisting of a QKV matrix and an MLP. CogVLM adopts a deep fusion approach, integrating visual and language features at multiple layers through the visual expert module, surpassing traditional shallow alignment methods. Alignment techniques involve pretraining on a vast dataset of 1.5 billion image-text pairs, employing image captioning loss and Referring Expression Comprehension (REC). Fine-tuning on various tasks, with a focus on free-form instructions, leads to the creation of a variant known as CogVLM-Chat.
@@
FERRET [104]: FERRET is designed for spatial referring and grounding in images at different shapes and granularities. Ferret’s distinct features include a Hybrid Region Representation, blending discrete coordinates and continuous visual features for diverse region inputs. It uses a Spatial-Aware Visual Sampler to handle various region shapes effectively and is trained on the Ground-and-Refer Instruction-Tuning (GRIT) dataset, which includes hierarchical spatial knowledge and hard negative samples. The architecture includes an image encoder, a spatial-aware visual sampler, and a Language Model . Ferret utilizes a pre-trained visual encoder (CLIP-ViT-L/14) and a Language Model’s tokenizer for image and text embeddings. Training occurs on the GRIT dataset for three epochs, with the model randomly choosing between center points or bounding boxes to represent regions. In multimodal chatting tasks, Ferret significantly enhances performance by integrating refer-and-ground capabilities. Notably, Ferret mitigates the issue of object hallucination, a common challenge in multimodal models.
@@
BARD[30]: BARD from Google utilizes a reinforcement learning framework to automate machine learning model design, architecture search, and hyperparameter tuning, making it accessible to users without extensive AI expertise. The system is positioned as a standalone experiment, with a focus on productivity, creativity, and curiosity enhancement. Users engage BARD for tasks such as writing resumes, creating workout routines, and planning itineraries. The model is pre-trained on diverse data sources, and responses are generated by considering context, classified against safety parameters, and re-ranked based on quality. Human feedback and evaluation, including fine-tuning and reinforcement learning on human feedback, are used to improve BARD. Limitations include potential inaccuracies, biases, persona attribution, false positives/negatives, and vulnerability to adversarial prompting. Google is committed to addressing these limitations and improving BARD responsibly over time.
@@
LLaMA-VID [44]: LLaMA-VID introduces a novel dual-token strategy, incorporating context and content tokens, to efficiently encode each video frame. This approach enables the model to handle hour-long videos while mitigating computational complexity. LLaMA-VID employs a hybrid architecture, incorporating pre-trained models like Vicuna for text processing and a Vision Transformer for image embeddings in videos. The Q-Former introduces the context-attention token (Et) by computing attention between query-generated text embeddings (Q) and visual tokens (X). Et encapsulates relevant visual features. A content token (Ev) is obtained through mean pooling on visual tokens. Both tokens are integrated into the V decoder for generating text responses. LLaMA-VID’s dual-token generation strategy, comprising context and content tokens, ensures adaptability to various settings, optimizing efficiency for videos while preserving detail for single images.LLaMA-VID is a video and image understanding model designed for efficiency, completing training in two days on 8xA100 GPUs. It uses EVA-G for visual encoding and QFormer for text decoding. The training set includes image and video caption pairs, with evaluations on diverse benchmarks. LLaMA-VID excels in zero-shot video QA benchmarks, achieving high accuracy with only two tokens per frame.
@@
CoVLM [43]: CoVLM introduces a novel approach to enhance large language models’ compositional reasoning by integrating vision-language communicative decoding. Utilizing communication tokens, the model dynamically composes visual entities and relationships, improving language generation through iterative communication with vision encoders and detection networks . CoVLM demonstrates strong performance across a range of tasks including visual reasoning, reading comprehension, and visual question answering. The model represents a noteworthy advancement in integrating vision and language models, with acknowledgment of potential future improvements in compositionality.
@@
Emu2: [75]: Emu2 stands as a generative multimodal model boasting 37 billion parameters, showcasing exceptional contextual learning across varied multimodal sequences. It sets unprecedented benchmarks in tasks requiring rapid comprehension with limited examples. Employing a unified autoregressive objective, Emu2 seamlessly integrates visual embeddings and textual tokens. Its architecture includes a visual encoder, multimodal modeling, and visual decoder, allowing coherent outputs across different modalities. Emu2 excels in vision-language tasks, instruction tuning, and controllable visual generation, showcasing state-of-the-art performance in image question answering, subject-driven generation, and zero-shot text-to-image generation. The paper acknowledges broader impact considerations and limitations, emphasizing responsible deployment in light of challenges such as hallucination, biases, and question-answering capabilities.
@@
Video-LLaMA [112]: Video-LLaMA is tailored to comprehend both the visual and auditory aspects of videos. By merging pre-trained visual and audio encoders with static LLMs, the model adeptly tackles the complexities of capturing temporal shifts in visual contexts while seamlessly integrating audio-visual cues. Utilizing a Video Q-former for temporal information and an Audio Q-former for audio encoding, the framework aligns audio-visual data with textual information. Experimental results demonstrate Video-LaMA’s effectiveness in comprehending video content and generating meaningful responses in audio and video-grounded conversations. However, the paper acknowledges limitations such as restricted perception capacities and challenges with long videos. Despite these, Video-LLaMA represents a notable advancement in audio-visual AI assistants, with the authors providing open-sourced resources for further development.
@@
Video-ChatGPT [60]: It is, a novel multimodal model enhancing video understanding by integrating a video-adapted visual encoder with a Large Language Model. The architecture leverages the CLIP ViT-L/14 visual encoder for spatiotemporal video representations and the V-v1.1 language model for comprehensive understanding. Notably, a dataset of 100,000 video-instruction pairs is created to fine-tune the model, focusing on temporal relationships and contextual understanding. The model exhibits competitive performance in correctness, detail orientation, contextual and temporal understanding, and consistency, surpassing contemporary models in zero-shot question-answering tasks. Qualitatively, Video-ChatGPT demonstrates proficiency in various video-based tasks but faces challenges in subtle temporal relationships and small visual details, indicating avenues for future improvement.
@@
LAVIN [58]: LAVIN is utilizing Mixture-of-Modality Adaptation (MMA) for cost-effective adaptation of LLMs to vision-language tasks. LaVIN, with lightweight adapters, achieves competitive performance and superior training efficiency in multimodal tasks like science question answering and dialogue. Remarkably, LaVIN requires only 1.4 training hours and 3.8M trainable parameters. Experimental results on the ScienceQA dataset demonstrate efficiency with comparable performance and reduced training time and storage costs. LAVIN represents a breakthrough in cost-effective adaptation but has limitations, including the potential for incorrect responses and challenges in identifying fine-grained details in images.
@@
BEiT-3 [86]: BEiT-3 represents a pioneering multimodal foundation model, embodying substantial integration across language, vision, and multimodal pretraining domains. Distinguished by its remarkable proficiency in transfer learning across both vision-centric and vision-language tasks, BEiT-3 underscores the advancement of convergence through innovative enhancements in backbone architecture, pretraining methodologies, and scalable model design. Leveraging Multiway Transformers, this model is characterized by a modular architecture facilitating profound fusion capabilities and modality-specific encoding. With a unified backbone, BEiT-3 executes cohesive masked ”language” modeling across images, English text, and image-text pairs, colloquially referred to as ”parallel sentences”. Empirical findings corroborate BEiT-3’s attainment of state-of-the-art performance benchmarks across a spectrum of tasks encompassing object detection, semantic segmentation, image classification, visual reasoning, visual question answering, image captioning, and cross-modal retrieval.
@@
mPLUG-2 [92]: mPLUG-2 leads the way by introducing a multi-module composition network, diverging from the traditional approach of sequence-to-sequence generation. This innovative design fosters modality collaboration while effectively addressing modality entanglement. The flexibility inherent in mPLUG-2 allows for the selective use of diverse modules across text, image, and video modalities for various understanding and generation tasks. Empirical evaluations demonstrate mPLUG-2’s prowess, achieving state-of-the-art or competitive results across an extensive range of over 30 downstream tasks. From challenging multi-modal endeavors like image-text and video-text understanding to uni-modal tasks spanning text-only, image-only, and video-only domains, mPLUG-2 exhibits its versatility. A notable achievement of mPLUG-2 is its groundbreaking performance, achieving a top-1 accuracy of 48.0 and an 80.3 CIDEr score in video QA and caption tasks. Impressively, these results were obtained with a significantly smaller model size and dataset scale. Furthermore, mPLUG-2 exhibits robust zero-shot transferability in both vision-language and video-language tasks, consolidating its forefront status in advancing multimodal pretraining methodologies.
@@
X2-VLM [108]: X2-VLM is a versatile model with a flexible modular architecture that integrates image-text and video-text pre-training into a unified framework. It excels in image-text and video-text tasks across different scales, balancing performance and model scale. X2-VLM’s modular design enhances transferability, allowing seamless use in various languages or domains. By substituting the text encoder, it outperforms state-of-the-art multilingual multimodal pre-trained models, demonstrating superior performance without requiring specific multilingual pre-training. This adaptability positions X2-VLM as a promising model in the field of multimodal pre-training.
Lyrics [56]: Lyrics introduces a novel approach to fine-tuning instruction and multimodal pretraining through cross-modal collaboration, built upon the foundational concepts of BLIP-2.It includes advanced techniques for refining visual inputs to extract specific visual characteristics, alongside modules designed for semantic segmentation, object detection, and image tagging. Within the Querying Transformer, visual features seamlessly merge with language inputs, enhanced by boundary boxes and tags derived from the visual refiner.A distinctive aspect of Lyrics is its two-stage training process, which aids in bridging the modality gap by aligning visual-language targets during pretraining. To extract valuable features from tangible objects, it employs a crucial technique called semantic-aware visual feature extraction. The effectiveness of this approach is evidenced by its robust performance across various visual-language benchmark tasks and datasets.
@@
X-FM [110]: XFM, a novel general foundation model equipped with one language encoder, one vision encoder, and one fusion encoder, featuring a unique training method. The proposed method incorporates two innovative techniques: halting gradients from vision-language training during language-encoder learning and leveraging vision-language training to guide vision-encoder learning. Extensive experiments on benchmark datasets demonstrate that X-FM outperforms existing general foundation models and performs competitively with or surpasses models tailored specifically for language, vision, or vision-language understanding. The paper acknowledges limitations, including substantial computational requirements, and aims to explore techniques for efficiency improvement and reduced environmental impact. The authors highlight their commitment to addressing efficiency challenges and reducing the carbon footprint in line with ”green” deep learning initiatives. However, due to computational constraints, the study did not explore super-large models or pre-train large-sized models on extensive datasets, emphasizing scalability as an essential consideration for foundation models.
@@
VALOR [11]: VALOR is a unified vision-audio-language crossmodality pretraining model designed for trimodality understanding and generation. VALOR employs two pretraining tasks, Multimodal Grouping Alignment and Multimodal Grouping Captioning, showcasing good versatility and scalability. Two datasets, namely VALOR-1M and VALOR-32K, emerge as pivotal resources for the advancement of tri-modality pretraining research, aimed at benchmarking audiovisual-language retrieval and captioning. Upon completion of training on the VALOR-1M dataset alongside other vision-language datasets, VALOR establishes novel performance benchmarks across diverse downstream tasks. These tasks encompass retrieval scenarios incorporating vision, audio, and audiovisual inputs, as well as tasks such as captioning and question answering. The documentation delineates prospective avenues for future research, notably including the expansion of the VALOR-1M dataset through unsupervised methodologies, alongside the integration of vision and audio generation modeling within the overarching VALOR framework.
@@
Prismer [52]: Prismer is a data- and parameter-efficient vision-language model that utilizes a frozen ensemble of domain experts, minimizing the need for extensive training data. By inheriting weights from pre-trained domain experts across various domains and keeping them frozen during training, Prismer efficiently adapts to different vision-language reasoning tasks. Despite its small-scale language model foundation, Prismer demonstrates competitive fine-tuned and few-shot learning performance, requiring significantly less training data than current state-of-the-art models. However, it lacks the ability for zero-shot in-context generalization and shows limitations in adapting to new experts or partial expert ensembles during inference, leading to performance drops. The paper discusses these limitations, including the absence of few-shot in-context prompting, challenges in adapting to new experts, and potential improvements in representing expert knowledge for enhanced reasoning performance in future iterations.
@@
MMReact [99]: MM-REACT introduces a novel textual prompt design enabling language models to process multimodal information, including text descriptions, spatial coordinates, and file names for dense visual signals. The approach demonstrates its effectiveness in zero-shot experiments, showcasing its potential for advanced visual understanding across various scenarios. However, the paper identifies limitations, such as challenges in systematically evaluating performance due to the absence of annotated benchmarks for recognition capability in the wild. The integrated vision experts may introduce errors, and the system’s success depends on the availability of necessary experts. Additionally, the number of experts is constrained by the context window of ChatGPT, and the conversion of visual signals to text words may not be optimal for certain tasks. Manual prompt engineering is required, and the authors suggest future research to automate this process for increased system development ease.
@@
PICa [37]: PICa is a method utilizing image captions to prompt GPT-3 for knowledge-based Visual Question Answering (VQA). Leveraging GPT-3’s knowledge retrieval and question-answering capabilities, the approach treats GPT-3 as an implicit and unstructured knowledge base, converting images into captions or tags for GPT-3 understanding. By adapting GPT-3 for VQA through a few-shot learning approach with in-context examples, PICa achieves notable performance, surpassing the supervised state of the art on the OK-VQA dataset with just 16 examples. The method is the first to use GPT-3 for multimodal tasks. However, a limitation is noted, as the image is abstracted as text, and captions may provide only a partial description, potentially missing crucial visual details necessary for detailed question answering, such as queries on specific visual attributes.
@@
PNP-VQA [80]: Plug-and-Play VQA (PNP-VQA) is a modular framework designed for zero-shot Visual Question Answering (VQA). Unlike existing approaches that demand extensive adaptation of pre-trained language models (PLMs) for vision, PNP-VQA eliminates the need for additional training of PLMs. Instead, it employs natural language and network interpretation as an intermediate representation to connect pre-trained models. The framework generates question-guided informative image captions, utilizing them as context for PLMs during question answering.PNP-VQA outperforms end-to-end trained baseline models and establishes new benchmarks by achieving state-of-the-art results on zero-shot VQAv2 and GQA datasets. Despite having 11 billion parameters, it surpasses the performance of an 80 billion-parameter model on VQAv2 and demonstrates a remarkable 9.1% improvement over a comparable model on GQA. This highlights its effectiveness across a range of parameter sizes for pre-trained language models (PLMs).
@@
Img2LLM [33]: Img2LLM is designed for LLMs that facilitates zero-shot VQA without necessitating end-to-end training. The approach involves developing LLM-agnostic models that articulate image content through exemplar question-answer pairs, proving to be effective prompts for LLMs. Img2LLM boasts several advantages, achieving performance on par with or surpassing end-to-end trained methods, such as outperforming Flamingo by 5.6% on VQAv2 and exhibiting notable superiority on the challenging A-OKVQA dataset. Additionally, the flexibility of Img2LLM allows seamless integration with various LLMs for VQA tasks, eliminating the need for specialized, costly end-to-end fine-tuning. One caveat is the additional inference overhead incurred during image caption and question-answer pair generation, contributing to a 24.4% increase in computational time. However, this overhead can be mitigated by shortening prompts, trading a fraction of accuracy for speed, while Img2LLM avoids the resource-intensive end-to-end multimodal representation alignment seen in comparable models like Flamingo.
@@
SimVLM [88]: SimVLM is a streamlined pretraining framework that embraces a minimalist approach. Unlike previous methods, SimVLM simplifies training complexities by leveraging large-scale weak supervision, and undergoing end-to-end training with a singular prefix language modeling objective. Remarkably, without resorting to additional data or task-specific tailoring, the resultant model surpasses its predecessors like OSCAR, VILLA etc establishing new benchmarks in various vision-language tasks. Additionally, SimVLM demonstrates robust generalization and transfer capabilities, showcasing zero-shot behavior in tasks such as open-ended visual question answering and cross-modality transfer.